# Measurement-only dynamical phase transition of topological and boundary orders in toric code and gauge-Higgs models

### v1

@@ Abstract
We extensively study long-time dynamics and fate of topologically-ordered state in toric code model evolving through a projective measurement-only circuit. The circuit is composed of several measurement operators corresponding to each term of toric code Hamiltonian with magnetic-field perturbations, which is a gauge-fixed version of a (2+1)-dimensional gauge-Higgs model. We employ a cylinder geometry to classify stationary states after long-time measurement dynamics. The appearing stationary states depend on measurement probabilities for each measurement operator. The Higgs, confined and deconfined phases exist in the time evolution by the circuit. We find that both the Higgs and confined phases are clearly separated from the deconfined phase by topological entanglement entropy, whereas the phase boundary between the Higgs and confined phases is obtained by long-range orders on the boundaries supporting the recent observation that the Higgs and confined phases are both one of symmetry-protected-topological states.
@@ I Introduction
Similar measurement-induced phase transitions are also observed in measurement-only circuit (MoC). With suitable choice of measurement operations and their frequencies, MoC generates non-trivial phases of matter; symmetry protected topological (SPT) phases Lavasani2021 ; Klocke2022 ; KI2023 , topological orders Lavasani2021_2 ; Negari2023 , and non-trivial thermal and critical phases Ippoliti2021 ; Sriram2023 ; KOI2023 ; Lavasani2023 ; Zhu2023 . These dynamical phenomena originate from non-commutativity and back action of measurements, and the interplay of them produces intriguing stationary states after ‘time evolution’ through circuit. Furthermore, suitable measurement on initially-prepared states can produce resource states for quantum computation Raussendorf2001 ; Briegel2009 , etc. Example of this ability includes production of long-range entanglement states with intrinsic topological order from simple symmetry-protected-topological (SPT) states Tantivasadakarn2022 .
A recent experiment Google_quantum_AI has realized toric code Kitaev2003 . The robustness of the system to decoherence error, measurements or magnetic perturbation is now open and one of the attracted issues Fan2023 ; Wang2023 .
@@
The above discussion was obtained by the Hamiltonian formalism. In this paper, we shall study circuit dynamics of toric code system with the cylinder geometry by using measurement-only dynamics. In particular, we investigate how initially-prepared toric code state (topological ordered state) dynamically changes. In this process, the choice of measurement operators and their probability are essential ingredients.
@@
We shall show that rich dynamical behavior and stationary states emerge in the system by varying protocol of the MoC. In particular, we find that Higgs and confined stationary states are separated by a critical line on the boundaries, and then they are explicitly distinguishable in the MoC, in contrast to the well-known conjecture in Fradkin1979 . Furthermore, when we apply strictly competitive measurement protocol for projective measurements, the saturation time gets very long, indicating that competitive dynamics takes place in critical regime among deconfined, Higgs and confined phases. On the other hand, we also observe that topological order (TO) of the initial toric code is sustained by frequent measurement of the toric code stabilizers to eliminate effects of the local projective measurements. This dynamical process can be useful for quantum error correction Botzung .
@@ II Measurement-only circuit and comparable Hamiltonian formalism
We shall investigate ‘time evolution’ of the pure stabilizer state specified by the following set of stabilizer generators Nielsen_Chuang , $S_{\rm int}=\{\tilde{A}_{v}|v\in\mbox{all }v\}+\{\tilde{B}_{q}|q\in\mbox{all }q\}$, where $\tilde{A}_{v}$ and $\tilde{B}_{q}$ are the star and plaquette operators of the toric code, defined by $\tilde{A}_{v}=\prod_{\ell_{v}\in v}\sigma^{x}_{\ell_{v}}$ and $\tilde{B}_{q}=\prod_{\ell_{q}\in q}\sigma^{z}_{\ell_{q}}$, with $\ell_{v}\in v$ standing for links emanating from vertex $v$, and $\ell_{q}\in q$ for links composing plaquette $q$. The above stabilizer generators are all linearly independent and specify the stabilizer state $S_{\rm int}$, which is nothing but the exact unique gapped ground state of the toric code Hamiltonian without magnetic perturbations on the cylinder geometry degene1 . Throughout this work, we use the same notation for the stabilizer set and the corresponding state interchangeably.
For the stabilizer state of this lattice system, we apply sequential projective measurements as MoC. In the protocol, we introduce two distinct layers: in the first one (called “layer 1”), two projective measurements with observables, $\hat{M}^{1}_{\ell}=X_{\ell}$ and $\hat{M}^{2}_{\ell}=Z_{\ell}$ are applied with a uniform probability $p_{x}$ and $p_{z}$, respectively, for each link except the bottom smooth boundary (in the case of $\hat{M}^{1}_{\ell}$) and each link except the top rough boundary (in the case of $\hat{M}^{2}_{\ell}$). We choose these measurement points inspired by the Hamiltonian we later commented on Eq. (1).
@@
The above model is a gauge-fixing version of a lattice gauge-Higgs model with open boundary conditions employed for investigation on its topological properties KOI2024 . The gauge-Higgs model on infinite system was analyzed in Fradkin and Shenker Fradkin1979 , where the phase diagram with Higgs, confined and deconfined(toric code) regimes was discovered. $H_{\rm TC}$ in Eq. (1) for $J_{x}=J_{z}=0$ with general boundary conditions is the fixed point Hamiltonian of the deconfined phase, and its ground states are a gapfull topological state called toric code. The model was recently re-investigated from viewpoint of SPT Verresen2022 , by employing specific boundary conditions. The Higgs and confined phases in (2+1) D can be distinguished with each other by observing the long-range order (LRO) on the boundaries, which is a characteristic signature implying that both the Higgs and confined phases are SPTs being protected by magnetic(electric)-one-form symmetry, respectively. For the Hamiltonian of Eq.(1), we expect a phase diagram such as: Higgs phase for $J_{z}>J_{x},h_{x},h_{z}$, confined phase for $J_{x}>J_{z},h_{x},h_{z}$ and deconfined(toric code) phase for $h_{x}=h_{z}>J_{z},J_{x}$ Trebst2007 ; Vidal2009 ; Tupitsyn2010 ; Dusuel2011 ; Wu2012 . The schematic image of the phase diagram is shown in Fig. 2.
Figure 2: Schematic phase diagram of the Hamiltonian $H_{\rm TC}$ in Eq. (1). Here, we fix $h_{x}=h_{z}$ to a finite value. By imposing a specific boundary conditions for a system with the cylinder geometry, the Higgs and confined phases are distinguished. SSB stands for spontaneous symmetry breaking.
@@
The above phase structure of the Hamiltonian of Eq. (1) gives us insights into the appearance of stationary states in the MoC started with the stabilizer state $S_{\rm int}$. In the previous study KI2023 , we investigated relationship between Hamiltonian systems and MoCs. The results obtained there indicate the following relationship between the parameter ratios in the present system, $|J_{x}|/|J_{z}|\longleftrightarrow p_{x}/p_{z}$ and $|h_{x(z)}|/|J_{x(z)}|\longleftrightarrow p_{s}/p_{x(z)}$.
@@
In the recent understanding of the above phase diagram of the gauge-Higgs model, which is an ancestor model of $H_{\rm TC}$ with exactly the same energy eigenstates, higher-form-symmetry plays an important role Gaiotto2015 ; McGreevy2022 . For the case of $J_{x}=0$ in $H_{\rm TC}$ in Eq. (1), there emerges magnetic-form-symmetry generated by the operator $W^{c}_{\gamma}\equiv\prod_{\ell\in\gamma}\sigma^{z}_{\ell}$, where $\gamma$ is an arbitrary closed loop on links of the $v$-lattice. Similarly for a string connecting two different links residing on the rough boundary, $\Gamma$, $W^{o}_{\Gamma}\equiv\prod_{\ell\in\Gamma}\sigma^{z}_{\ell}$ is another one-form-symmetry. These one-form-symmetries can be robust for a finite value of $J_{x}$ as long as the state belongs to the same phase with that of $J_{x}=0$. More precisely, the deconfined(toric code) phase is characterized as an SSB phase of the $W^{c}_{\gamma}$ and $W^{o}_{\Gamma}$ symmetries. On the other hand, the Higgs phase is an SPT of the above two symmetries Verresen2022 . For $J_{z}=0$, parallel discussion works for the confined phase as a duality picture and the electric-form-symmetries, $H^{c}_{\tilde{\gamma}}\equiv\prod_{\ell\in\tilde{\gamma}}\sigma^{x}_{\ell}$ and $H^{o}_{\tilde{\Gamma}}\equiv\prod_{\ell\in\tilde{\Gamma}}\sigma^{x}_{\ell}$, where links $\ell$’s are crossed by $\tilde{\gamma}$ and $\tilde{\Gamma}$, a close loop and string (connecting two links on the smooth boundary) on the dual lattice, respectively.
@@ III Numerical stabilizer simulation
Figure 3: (i) Long-time dynamics of Edward-Anderson spin-glass (SG) order parameter for $p_{s}=0.5$ and various values of $p_{x}=1-p_{z}$. The SG order exhibits different behavior depending on $p_{x}$. (ii) Sample-to-sample variance for the SG order. (iii) The late time saturation value of the SG order. We took both sample and time average. As for the time average, we employ the last nine data points. We used $1000$, $1000$, and $500$ samples for $L_{x}=14,18,$ and $22$.
@@
This work clarifies the following problems in the MoC; (i) how and when boundary LROs emerge on the upper rough and lower smooth boundaries (ii) how and when the bulk TO disappears. In the numerical study addressing the above problems, we fix $L_{y}=10$ and vary $L_{x}$ up to 26. For the first problem, we calculate the Edward-Anderson SG order parameter Sang2021 defined by
@@
with $C_{SG}(\ell_{x},\ell^{\prime}_{x})=\langle\psi|\sigma^{z}_{\ell_{x}}\sigma^{z}_{\ell^{\prime}_{x}}|\psi\rangle^{2}-\langle\psi|\sigma^{z}_{\ell_{x}}|\psi\rangle^{2}\langle\psi|\sigma^{z}_{\ell^{\prime}_{x}}|\psi\rangle^{2}$, where $|\psi\rangle$ is a stabilizer state and $\ell_{x}$’s are links on the boundary. As a target observable, we observe the variance of $\chi_{SG}$ divided by $L_{x}$, $F_{v}(p)\equiv\mbox{var}(\chi_{SG})/L_{x}$. This quantity measures sample-to-sample fluctuations and is useful to search a phase transition on the boundary. For the second problem, we calculate topological entanglement entropy (TEE), $\gamma$. In the present study, we employ the partition of the system shown in Fig. 4 Kitaev_Preskill , and for that partition, $\gamma$ is defined as
@@
where $S_{A},S_{B},\cdots$ denote entanglement entropy of the corresponding subsystem, which can be calculated from the number of linearly-independent stabilizers within a target subsystem and the number of qubit of the subsystem Fattal2004 ; Nahum2017 . If the system is an exact topologically-ordered state, we have $\gamma=-1$. The TEE is a useful and standard quantity for investigating bulk TO in the MoC. In particular, even for one of topological degenerate ground states, the TEE works efficiently to observe the TO. In addition, we consider another partition for TEE Levin_Wen , the numerical results of which are given in Appendix B.

### v2

# Measurement-only dynamical phase transition of topological and boundary order in toric code and gauge-Higgs models

@@ Abstract
We extensively study long-time dynamics and fate of topologically-ordered state in toric code model evolving through projective measurement-only circuit. The circuit is composed of several measurement operators corresponding to each term of toric code Hamiltonian with magnetic-field perturbations, which is a gauge-fixed version of a (2+1)-dimensional gauge-Higgs model. We employ a cylinder geometry with distinct upper and lower boundaries to classify stationary states after long-time measurement dynamics. The appearing stationary states depend on measurement probabilities for each measurement operator. The Higgs, confined and deconfined phases emerge in the time evolution by the circuit. We find that both Higgs and confined phases are separated from the deconfined phase by topological entanglement entropy. We numerically clarify that both Higgs and confined phases are characterized by a long-range order on the boundaries accompanying edge modes, and they shift with each other by a crossover reflecting properties in the bulk phase diagram.
@@ I Introduction
Similar measurement-induced phase transitions are also observed in measurement-only circuit (MoC). With suitable choice of measurement operations and their frequencies, MoC generates non-trivial phases of matter; symmetry protected topological (SPT) phases Lavasani2021 ; Klocke2022 ; KI2023 , topological orders Lavasani2021_2 ; Negari2023 , and non-trivial thermal and critical phases Ippoliti2021 ; Sriram2023 ; KOI2023 ; Lavasani2023 ; Zhu2023 . These dynamical phenomena originate from non-commutativity and back action of measurements, and the interplay of them produces intriguing stationary states after ‘time evolution’ through circuit. Furthermore, suitable measurement on initially-prepared states can produce resource states for quantum computation Raussendorf2001 ; Briegel2009 , etc. Example of this ability includes production of long-range entanglement states with intrinsic topological order from simple SPT states Tantivasadakarn2022 .
A recent experiment Google_quantum_AI has realized toric code Kitaev2003 . The robustness of the system to decoherence error, measurements or magnetic perturbation is now open and one of the attracted issues J-Y-Lee2023 ; Fan2023 ; Wang2023 .
@@
The above discussion was obtained by the Hamiltonian formalism. In this paper, we shall study circuit dynamics of toric code system with the cylinder geometry by using measurement-only dynamics, where we employ distinct upper and lower boundaries, elucidating a duality between the Higgs and confined phases. This gives an efficient setup for further investigation of the behavior of the transition between the Higgs and confined phases. In particular, we study how initially-prepared toric code state (topological ordered state) dynamically changes. In this process, the choice of measurement operators and their probability are essential ingredients.
@@
We shall show that rich dynamical behavior and stationary states emerge in the system by varying protocol of the MoC. In particular, we find that Higgs and confined stationary states are separated via crossover behavior under this geometry, and then they are observed in the MoC by the boundary orders, in contrast to the well-known observation in Fradkin1979 . According to our previous work KI2023_2 , we expect that the obtained result gives an important insight into the boundary physics of the corresponding Hamiltonian system, for which only qualitative conjecture has been given so far. Furthermore, when we apply strictly competitive protocol for projective measurements, the saturation time gets very long, indicating that competitive dynamics takes place in critical regime among deconfined, Higgs and confined phases. On the other hand, we also observe that topological order (TO) of the initial toric code under this specific boundary conditions is sustained by frequent measurement of the toric code stabilizers as a result of reduction of the projective-measurement effects. This dynamical process can be useful for quantum error correction Botzung .
@@ II Measurement-only circuit and comparable Hamiltonian formalism
We shall investigate ‘time evolution’ of the pure stabilizer state specified by the following set of stabilizer generators Nielsen_Chuang , $S_{\rm int}=\{\tilde{A}_{v}|v\in\mbox{all }v\}+\{\tilde{B}_{q}|q\in\mbox{all }q\}$, where $\tilde{A}_{v}$ and $\tilde{B}_{q}$ are the star and plaquette operators of the toric code, defined by $\tilde{A}_{v}=\prod_{\ell_{v}\in v}\sigma^{x}_{\ell_{v}}$ and $\tilde{B}_{q}=\prod_{\ell_{q}\in q}\sigma^{z}_{\ell_{q}}$, with $\ell_{v}\in v$ standing for links emanating from vertex $v$, and $\ell_{q}\in q$ for links composing plaquette $q$. The above stabilizer generators are all linearly independent and specify the stabilizer state $S_{\rm int}$, which is nothing but the exact unique gapped ground state of the toric code Hamiltonian without magnetic perturbations on the cylinder geometry degene1 . The motivation to select this geometry and unique pure states as an initial state is that we can observe some duality between the upper rough and lower smooth boundary that can elucidate the Higgs and confined phase, while we expect the bulk physics of stationary states is not significantly affected by the geometry and choice of initial states. Throughout this work, we use the same notation for the stabilizer set and the corresponding state interchangeably.
For the stabilizer state of this lattice system, we apply sequential projective measurements as MoC. In the protocol, we introduce two distinct layers: in the first one (called “layer 1”), two projective measurements with observables, $\hat{M}^{1}_{\ell}=\sigma^{x}_{\ell}$ and $\hat{M}^{2}_{\ell}=\sigma^{z}_{\ell}$ are applied with a uniform probability $p_{x}$ and $p_{z}$, respectively, for each link except the bottom smooth boundary (in the case of $\hat{M}^{1}_{\ell}$) and each link except the top rough boundary (in the case of $\hat{M}^{2}_{\ell}$). We choose these measurement points inspired by the Hamiltonian we later commented on Eq. (1).
@@
The above model is a gauge-fixing version of a lattice gauge-Higgs model with open boundary conditions employed for investigation on its topological properties KOI2024 . In this work, we assume $h_{x(z)}$, $J_{x(z)}\geq 0$. The gauge-Higgs model on infinite system was analyzed in Fradkin and Shenker Fradkin1979 , where the phase diagram with Higgs, confined and deconfined (toric code) regimes was discovered. $H_{\rm TC}$ in Eq. (1) for $J_{x}=J_{z}=0$ with general boundary conditions is the fixed point Hamiltonian of the deconfined phase, and its ground states are a gapped topological state called toric code. Gauge-Higgs model related to the above TC model was recently re-investigated from viewpoint of SPT Verresen2022 , by employing specific boundary conditions. In this work, we extend the above study by introducing the cylinder geometry with both rough and smooth boundaries as in Eq. (1). Then, the Higgs and confined phases in (2+1) D can be distinguished with each other by observing the long-range order (LRO) on the boundaries, which is a characteristic signature implying that both the Higgs and confined phases are SPTs being protected by magnetic(electric)-one-form symmetry, respectively. Findings in this present work are, therefore, complementary to the study in Ref. Verresen2022 .
It is verified by the practical calculation KI2023_2 that the gauge-Higgs and TC models are exactly connected by the gauge fixing (so called unitary gauge) even with the specific open boundary conditions employed in Ref. Verresen2022 and also in the present study. This is quite natural as local gauge transformation can be performed at each site where the matter field resides on. (For more detailed discussion, in particular for the bulk-edge entanglement, see Sec. III F of Ref. KI2023_2 .) Therefore for the Hamiltonian of Eq.(1), we expect a phase diagram such as: Higgs phase for $J_{z}>J_{x},h_{x},h_{z}$, confined phase for $J_{x}>J_{z},h_{x},h_{z}$ and deconfined (toric code) phase for $h_{x}=h_{z}>J_{z},J_{x}$ Trebst2007 ; Vidal2009 ; Tupitsyn2010 ; Dusuel2011 ; Wu2012 . The schematic image of the phase diagram is shown in Fig. 2, where the schematic is based on the global phase structure of the TC model with magnetic fields under periodic boundary, which has been clarified by using Monte Carlo numerical calculation Tupitsyn2010 . This phase diagram of the Hamiltonian of Eq. (1) gives us insights into the appearance of stationary states in the MoC started with the stabilizer state $S_{\rm int}$.
@@
In the previous study KI2023 , we investigated relationship between Hamiltonian systems and MoCs. A brief explanation is shown in Appendix A. The results obtained there indicate the following relationship between the parameter ratios in the present system, $J_{x}/J_{z}\longleftrightarrow p_{x}/p_{z}$ and $h_{x(z)}/J_{x(z)}\longleftrightarrow p_{s}/p_{x(z)}$.
Figure 2: Schematic phase diagram of the Hamiltonian $H_{\rm TC}$ in Eq. (1). Here, we fix $h_{x}=h_{z}$ to a finite value. By imposing a specific boundary conditions for a system with the cylinder geometry, the Higgs and confined phases are distinguished. The label “SSB” stands for spontaneous symmetry breaking. We expect the relationship between the parameter ratio and the probability ratio, $J_{x}/J_{z}\longleftrightarrow p_{x}/p_{z}$ and $h_{x(z)}/J_{x(z)}\longleftrightarrow p_{s}/p_{x(z)}$, which is a good guide to expect the steady state in the MoC. The blue dotted line is expected to be $p_{x}=p_{z}$ case in our MoC. Double-headed arrows (i) and (ii) represent the parameter spaces we focus on in this study. In particular, we focus on (i) in Sec. IV and (ii) in Sec. V.
@@
In the recent understanding of the above phase diagram of the gauge-Higgs model, which is an ancestor model of $H_{\rm TC}$ with exactly the same energy eigenstates, higher-form-symmetry plays an important role Gaiotto2015 ; McGreevy2022 . For the case of $J_{x}=0$ in $H_{\rm TC}$ in Eq. (1), there emerges magnetic-form-symmetry generated by the operator $W^{c}_{\gamma}\equiv\prod_{\ell\in\gamma}\sigma^{z}_{\ell}$, where $\gamma$ is an arbitrary closed loop on links of the $v$-lattice. Similarly for a string connecting two different links residing on the rough boundary, $\Gamma$, $W^{o}_{\Gamma}\equiv\prod_{\ell\in\Gamma}\sigma^{z}_{\ell}$ is another one-form-symmetry (See Appendix B). These one-form-symmetries can be robust for a finite value of $J_{x}$ as long as the state belongs to the same phase with that of $J_{x}=0$. More precisely, the deconfined (toric code) phase is characterized as an SSB phase of the $W^{c}_{\gamma}$ and $W^{o}_{\Gamma}$ symmetries. On the other hand, the Higgs phase is an SPT of the above two symmetries Verresen2022 . For $J_{z}=0$, parallel discussion works for the confined phase as a duality picture and the electric-form-symmetries, $H^{c}_{\tilde{\gamma}}\equiv\prod_{\ell\in\tilde{\gamma}}\sigma^{x}_{\ell}$ and $H^{o}_{\tilde{\Gamma}}\equiv\prod_{\ell\in\tilde{\Gamma}}\sigma^{x}_{\ell}$, where links $\ell$’s are crossed by $\tilde{\gamma}$ and $\tilde{\Gamma}$, a close loop and string (connecting two links on the smooth boundary) on the dual lattice, respectively.
An interesting issue is whether the SPT property of the Higgs phase in the gauge-Higgs model is preserved in the TC model obtained from the gauge-Higgs model through the gauge fixing. An important observation concerning this issue is that ‘logical operators’ such as $W^{o}_{\Gamma}\equiv\prod_{\ell\in\Gamma}\sigma^{z}_{\ell}$ with $\Gamma$ embedded in the smooth boundary, which play an essentially important role for edge-mode physics on the boundaries, are invariant under the gauge fixing as well as related unitary transformation KI2023_2 ; Wildeboer2022 . This is quite natural as these are gauge-invariant operators in the gauge-Higgs model. The SSB of symmetries produced by the ‘logical operators’ is evidence of the existence of SPT, which we verify in the present study. In other words, one-form symmetries in the gauge-Higgs and TC models play essentially the same role for the SPT properties. Simply put, it is plausible to expect that Higgs=SPT is a gauge-invariant property if it exists.
Finally, an important remark concerning the phase diagram in Fig. 2 is in order. How the boundary physics looks like in the Higgs-confined ‘critical regime’ is an interesting unsolved problem, even though a schematic line separating the Higgs and confined phases is drawn in a phase diagram in Ref. Verresen2022 . In the previous work studying the boundary physics in the critical regime of the deconfined-Higgs phase transition KOI2024 , we showed that a genuine transition-like behavior emerges on the boundary as a result of the strict bulk phase transition. On the other hand as the Higgs and confined phases are connected in an adiabatic matter without any singular behaviors, it is important to investigate the boundary physics in that regime. We shall study this problem in this work to obtain an interesting result. We expect that the result obtained by the present MoC gives an important insight into the ground-state phase diagram of the corresponding Hamiltonian system KI2023 .
@@ III Numerical stabilizer simulation
Figure 3: (i) Long-time dynamics of Edward-Anderson spin-glass (SG) order parameter for $p_{s}=0.5$ and various values of $p_{x}=1-p_{z}$. The SG order exhibits different behavior depending on $p_{x}$. (ii) Sample-to-sample variance for the SG order. (iii) The late time saturation value of the SG order. The inset of (iii) shows $\sigma(\chi_{SG_{x(z)}})/L_{x}$, and we used the same data as in (iii). We took both sample and time average. As for the time average, we employ the last nine data points. We used $2000$, $2000$, and $1000$ samples for $L_{x}=14,18$ and $22$, respectively.
@@
This work clarifies the following problems in the MoC; (i) how and when boundary LROs emerge on the upper rough and lower smooth boundaries (ii) how and when the bulk TO disappears. In the numerical study addressing the above problems, we fix $L_{y}=10$ and vary $L_{x}$ up to 26. For the first problem, one may expect that realizing LROs can be identified by a standard long-rand order parameter, e.g., $\langle\psi|\sigma^{z(x)}_{\ell_{x}}\sigma^{z(x)}_{\ell^{\prime}_{x}}|\psi\rangle-\langle\psi|\sigma^{z(x)}_{\ell_{x}}|\psi\rangle\langle\psi|\sigma^{z(x)}_{\ell^{\prime}_{x}}|\psi\rangle$; however, this is not the case. This is because realized LROs due to measurement are more likely to be a glassy-type LRO, which closely relates to each measurement outcome. The glassy-type LRO is to be measured by an order parameter that is independent of direction of outcome ‘spins’ and merely depends on the magnitude of their correlations. Thus, to evaluate the glassy type LRO, we employ the Edward-Anderson SG order parameter Sang2021 defined by
@@
with $C_{SG}(\ell_{x},\ell^{\prime}_{x})=\langle\psi|\sigma^{z}_{\ell_{x}}\sigma^{z}_{\ell^{\prime}_{x}}|\psi\rangle^{2}-\langle\psi|\sigma^{z}_{\ell_{x}}|\psi\rangle^{2}\langle\psi|\sigma^{z}_{\ell^{\prime}_{x}}|\psi\rangle^{2}$, where $|\psi\rangle$ is a stabilizer state and $\ell_{x}$’s are links on the boundary. As a target observable, we observe the variance of $\chi_{SG}$ divided by $L_{x}$, $\sigma(\chi_{SG})/L_{x}$. This quantity measures sample-to-sample fluctuations and is useful to search a phase transition on the boundary. For the second problem, we calculate topological entanglement entropy (TEE), $\gamma$. In the present study, we employ the partition of the system shown in Fig. 4 Kitaev_Preskill , and for that partition, $\gamma$ is defined as
@@
where $S_{A},S_{B},\cdots$ denote entanglement entropy of the corresponding subsystem, which can be calculated from the number of linearly-independent stabilizers within a target subsystem and the number of qubit of the subsystem Fattal2004 ; Nahum2017 . The technical details of the methods to calculate EE are given in Appendix A in Ref. KOI2023 and Appendix E in Ref. Gullans2020 . Here, we employ EE with base-2 logarithm, following to the conventional definition of quantum information. As one example, the EE for $S_{A}$ is given by $S_{A}=-\operatorname{tr}[\rho_{A}\log_{2}\rho_{A}]$ where $\rho_{A}$ is reduced density matrix for A-subsystem. If the system is an exact topologically-ordered state, we have $\gamma=-1$. The TEE is a useful and standard quantity for investigating bulk TO in the MoC. In particular, even for one of topological degenerate ground states, the TEE works efficiently to observe the TO. In addition, we consider another partition for TEE Levin_Wen , the numerical results of which are given in Appendix C.